Prev Next

SW&IT Meeting Minutes 2018-08-13

Attendees: Andrew, David, Igor, Matthieu, Petr, Xianyu, Yang


News from management

  • candidate on-site interview Wed, SWG session @ 12-1pm, light lunch + zoom
  • Board's request for an AO Plan
  • Argos lead

Previous Action Items

  • Vsphere (DE)
  • IT email format (DE)
  • LBTI/jumbo switch (DE)

Open topics

Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.

  • trivial passwords (ovms-uei, others?) (SH when he comes back)
  • wiki software cleanup (MB, mainly due to his own contribution: week reports, etc. - but others too)

This week in github

Team reports

<!-- Enter your own using format:

---+++ <initials>



  • Work on server room power plan update
  • OVMS SD card replacement finally seems to work
  • Moved obs room Polycom to large monitor
  • Switched to internal IT email list for ticket type Instrument. Little activity, but what has gone out looks positive.
  • Coordinating with David Carroll on ticket type safety
  • Work on AGw computers
    • Replace old spinning disks with SSDs
    • Remove and clean controller cards/riser card for ideal electrical contact
    • Disable Windows startup check for smoother headless operation
    • Ordered spare riser cards
  • Started quote process for additional DIN-mount network switch (will follow up today)
  • Recieved Cisco swtich + fiber SFP modules to support jumbo frames
  • Research into VMWare upgrade
This week:
  • VMWare upgrade
  • Server room once-over to look for any PM tasks we can do during this shutdown
  • Reminder: MGIO power cut starting Tuesday 0800 through ~EOD Wednesday
  • Power upgrade plan


  • Continue on Python work:
    • Finishing up and improving collimation, fitting, and data searching utilities
    • Experimenting on Python documentation, sphinx, Python notebook, etc.
    • Work on other related documentation on pointing model.
  • Continue to study the AO supervisor code and the IT issue #7261 (memory problem with WfsArbitrator).


  • management OC, (no WPC)
  • interview 2 more candidates, final short list, interview schedule
  • labor reporting model cleanup (active staff, groups)
  • pp2018 completion, pp2019 planning
  • SOUL jumbo frame planning for LBTI
  • SHARK-NIR iteration with Marco
  • IQ/AO tool discussion with Al
  • aluminizing stop-motion movie (having fun with analysis)
  • candidate interview Wed.
  • pp2018 completion, pp2019 planning
  • IT to github issues
  • Query OVMS racktangle status with James/Leroy
  • eng. / attic repos move to git
  • TCS release planning meet with PK on Aug. 20
  • SOUL jumbo frame planning for LBTI
  • SHARK-NIR iteration with Marco


Last week
  • AGW3 in the lab needed the azcamserver software change done on AGW4 for them to do testing. I updated the software, but it developed a hardware issue recognizing the boards. Updated notes on the wiki for the software change that sets up sequential filenames. (IT7355)
  • OVMS kernel update by Petr/LeRoy/Stephen is working for 11 boards, but we get the same error as originally (Nov-2017) when trying to run 12 boards. Petr thought we should be using the 217 API, but our model is 211, so we're using the right API. I went back through all the discussion when we first saw this problem, and used some of the example programs to get some info on what the UEI thinks.
  • Committed the last couple 2018A GCS updates for OAC HK modifications - nicer logging of OAC errors.
  • Checked 2018B GCS DD entries getting set correctly from OAC. Waiting for Petr to merge 2018A to master and then I will update GCS in master.
  • Checked in the OAC updates from the last couple weeks, including some updates this week: check status of setthreshold commands; fixed my agw/pmac number terminology; updated config file with the mountain version. Built this version from SVN on the mountain as 5.0.

This week:
  • vacation


Last week:
  • Python script to remove headers as links on ARGOS wiki
  • Updated SVN links to point to GitHub
  • ARGOS deployment
  • Mocked ideas for ARGOS on FACSUM

This week
  • Vacation on Friday
  • ARGOS swing arm (see PK)
  • Deploy python script to fix ARGOS wiki
  • Continue ARGOS deployment on ops machines


  • More work on ARGOS, decided to introduce AbstractArm as the current software counts with two arm types already (OSS SwingArms and future Mirror cover; ARGOS will be the third type), code looks fine and compiles, need to implement Argos (and possible split what is done for MirrorCover intro separate class).

  • "Vacation" (up to Thursday China trip, presenting LBTO on Tuesday at Yunnan observatory; friend's wedding on Friday, hopefully some fun over weekend as we join Prague's Observatory field trip/expeditions and see some Perseids)
  • waiting for somebody to approve PR28, so I can merge it to 2018A and start pull request from 2018A back to master to start (sometime soon) 2018B


Last week:
  • Follow DX AdSec lab work;
  • Write AdSec status machine transitions with power backplance;
  • Check the AdSec electrinics calibrattion configration file with different shells;
  • Report AdSec SW/Safety with SOUL
This week:
  • Write the functionalities for AdSec status machine with power backplance?
  • support some other summer shutdown activities
  • read the AdSec electrinics calibrattion paper
  • FLAO SOUL WFS lab test


This week:
  • LUCI work:
    • FR#6 Adding AFC top and bottom pinhole selection to MOS tab in LUCI RTD
    • Work on LUCI handover document
  • PIT work:
    • Take out any italian logic of the general PIT
    • Possibly add more concise table on PDF
    • Other stuff partners might require, still TBD
Next week

Depending on LUCI availability and testing schedule. If LUCI is ready for testing with Dave, probably we'll be debugging and supporting for AFC work.

Topic revision: r17 - 10 Dec 2018, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback